OpenStudy (anonymous):

find the real solution 2x+6 -x+10=1

OpenStudy (anonymous):

\[\sqrt{2x+6}-\sqrt{x+10}=1\]

OpenStudy (anonymous):

help

OpenStudy (anonymous):

can someone help me please